# Distributed Motion Control of Multiple Mobile Manipulators for Reducing Interaction Wrench in Object Manipulation

### v2

@@ III-A Problem Statement and Control Law
Assumption 3: The motion of robots in cooperative object manipulation is a quasi-static process.
Assumption 1 is a very general assumption, which can also be found in [21, 11, 31, 32, 33]. Other assumptions are further discussed in subsequent theoretical proofs and experiments.
For the planning of cooperative manipulation of objects by multiple mobile manipulator robots, there are several references. This paper considers the motion control problem of executing a planned cooperative manipulation trajectory. For the sake of brevity, $\bm{q}(t)$ is abbreviated as $\bm{q}$ in the absence of time delay. For each robot, assuming that $\bm{h}_{i}^{\text{ref}}$, $\dot{\bm{h}}_{i}^{\text{ref}}$, $\bm{q}_{\text{mm},i}^{\text{ref}}$, and $\dot{\bm{q}}_{\text{mm},i}^{\text{ref}}$ have been planned. Since the manipulation is a quasi-static process, the desired wrench of the object $\bm{h}_{o}^{d}$ equals the gravity of the object. Then, $\bm{h}_{i}^{\text{ref}}$ and $\dot{\bm{h}}_{i}^{\text{ref}}$ can be calculated by
$\left[\begin{matrix}\bm{h}_{1}^{\text{T}}&\cdots&\bm{h}_{n}^{\text{T}}\end{matrix}\right]^{\text{T}}=-\bm{G}^{\dagger}\bm{h}_{o}^{d}$ (6)
where $\bm{G}^{\dagger}$ is the generalized inverse of $\bm{G}$. The load distribution $\bm{h}_{i}^{\text{ref}}$ is achieved based on the specific form of $\bm{G}^{\dagger}$ [30]. The motions are planned according to the desired trajectory of the object.
The desired trajectories and interaction wrenches are achieved if $\dot{\bm{q}}_{\text{mm},i}^{\text{ref}}$ is the input. However, due to model deviations and disturbances, this ideal situation is almost impossible. If global positioning and velocity information could be obtained, the correction process would be relatively easy since the actual velocity is corrected to match the desired velocity. However, this often requires high-precision and large-scale positioning systems, which are often difficult to implement [26]. Therefore, information from local force sensors is considered to design the control laws. Besides, it is also important to consider the communication delay between robots as it is a very common situation in reality. Let $\tau_{ij}(t)$ denote the time delay in the $i$-th robot receiving information from the $j$-th robot. In this paper, the most general delays are considered. The communication delays are non-uniform, asymmetric, time-varying, and unknown. For example, $\tau_{12}(t)\neq\tau_{21}(t)$, $\tau_{12}(t)\neq\tau_{13}(t)$, and thus robots cannot achieve synchronization by querying their own historical information when receiving unknown communication delays.
Since the object is not an absolute rigid body, the actual force exerted on the $i$-th robot during motion by the object is
@@
where $K_{6\times 6}>0$ is a diagonal matrix, representing the stiffness in six dimensions of the object. Eq.(7) can be rewritten as
@@
where $\bm{e}_{i}^{\text{d}}$ is the desired correction target. Through the force sensor, the direction of correction can be obtained.
It is well known that the manipulator is more accurate and controllable than the mobile platform [34]. In general, deviations in the executed planned trajectory are relatively small and can be corrected within the workspace of the manipulator. Therefore, the desired correction target is achieved through the manipulator. The control law of the manipulator for maintaining desired interaction wrenches is as follows
$\bm{u}_{i}=-k$ $\sum_{j=1}^{n}a_{ij}\Big{[}f^{-1}(K_{0}^{-1}(\bm{h}_{i}^{\text{ref}}-\bm{h}_{i}))$ (9)
$-\beta f^{-1}(K_{0}^{-1}(\bm{h}_{j}^{\text{ref}}(t-\tau_{ij})-\bm{h}_{j}(t-\tau_{ij})))\Big{]}$
where $k$ and $\beta$ are positive coefficients, $K_{0}$ is an estimate or empirical value of the stiffness, and $f^{-1}(\cdot)$ is the inverse kinematics of the manipulator. The input required for this control law is the local force sensor information, current joint position of the manipulator, and the information from neighboring robots. The control law ensures that the end of the robot moves towards the desired target $\bm{e}_{i}\to\bm{e}_{i}^{\text{d}}$, thereby maintaining the interaction wrench at the expected value $\bm{h}_{i}\to\bm{h}_{i}^{\text{ref}}$.
@@
where $\bm{q}_{i}^{\text{d}}$ is the desired correction joint target. Substituting Eq.(10) into Eq.(9) and let $\kappa=kK^{-1}_{0}K$, the control law can be rewritten as
$\bm{u}_{i}=-\kappa\sum_{j=1}^{n}{a_{ij}((\bm{q}_{i}-\bm{q}_{i}^{\text{d}})-\beta(\bm{q}_{j}(t-\tau_{ij})-\bm{q}_{j}^{\text{d}}(t-\tau_{ij})))}$ (11)
Eq.(9) is the control law used in actual calculations, indicating the input and output. The control rate is written as Eq.(11) to conveniently illustrate subsequent proofs of stability.
III-B Stability Analysis
@@
The joint position error of the $i$-th robot is
@@
$\dot{\bm{\xi}}_{i}=-\kappa\sum_{j=1}^{n}{a_{ij}}{\bm{\xi}}_{i}+\kappa\beta\sum_{j=1}^{n}{a_{ij}\bm{\xi}_{j}(t-\tau_{ij})}$ (17)
For the sake of brevity, the quadratic type $\bm{\xi}_{i}^{\text{T}}\kappa\bm{\xi}_{i}$ is abbreviated as $\kappa\bm{\xi}_{i}^{2}$. $I_{n}$ denotes the n-dimensional identity matrix. Considering the following Lyapunov function
$\quad\quad\quad\quad V(t)=V_{1}+V_{2}+V_{3}$ (18)
@@
where $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ are the positive coefficients of the Lyapunov function. Taking the time derivative of the Lyapunov function obtains
@@
Then, each part of Eq.(19) is scaled. According to Assumption 2, substituting Eq.(17) into $\dot{V}_{1}$ obtains
$\dot{V}_{1}$ $=-2\gamma_{1}\kappa\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}\bm{\xi}_{i}^{2}+2\gamma_{1}\kappa\beta\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}\bm{\xi}_{i}^{\text{T}}\bm{\xi}_{j}(t-\tau_{ij})$ (20)
$\leq-2\gamma_{1}\kappa\sum_{i=1}^{n}\bm{\xi}_{i}^{2}+2\gamma_{1}\beta\kappa\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}\bm{\xi}_{i}^{\text{T}}\bm{\xi}_{j}(t-\tau_{ij})$
@@
$\dot{V}_{2}$ $=\gamma_{2}\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ji}\bm{\xi}_{i}^{2}-\gamma_{2}\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ji}\bm{\xi}_{i}^{2}(t-\tau)$ (21)
$\leq\gamma_{2}\sum_{i=1}^{n}n\bm{\xi}_{i}^{2}-\gamma_{2}\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ji}\bm{\xi}_{i}^{2}(t-\tau)$
@@
$\int_{t-\tau}^{t}{\dot{\bm{\xi}}_{i}^{2}(s)}\text{d}s\geq\frac{(\bm{\xi}_{i}(t-\tau_{ji})-\bm{\xi}_{i}(t-\tau))^{2}}{\tau}$ (24)
$\quad+\frac{(\bm{\xi}_{i}-\bm{\xi}_{i}(t-\tau_{ji}))^{2}}{\tau}$
$\quad+\frac{(\bm{\xi}_{i}(t-\tau_{ji})-\bm{\xi}_{i}(t-\tau))^{\text{T}}(\bm{\xi}_{i}-\bm{\xi}_{i}(t-\tau_{ji}))}{\tau}$
$=\frac{1}{\tau}\Big{[}\bm{\xi}_{i}^{2}(t-\tau_{ji})+\bm{\xi}_{i}^{2}(t-\tau)+\bm{\xi}_{i}^{2}-\bm{\xi}_{i}^{\text{T}}\bm{\xi}_{i}(t-\tau)$
$\quad-\bm{\xi}_{i}^{\text{T}}\bm{\xi}_{i}(t-\tau_{ji})-\bm{\xi}_{i}^{\text{T}}(t-\tau_{ji})\bm{\xi}_{i}(t-\tau)\Big{]}$
$=\frac{1}{\tau}\Big{[}\frac{1}{2}\bm{\xi}_{i}^{2}(t-\tau_{ji})+(\frac{1}{2}\bm{\xi}_{i}(t-\tau_{ji})-\bm{\xi}_{i}(t-\tau))^{2}$
$\quad+(\frac{1}{2}\bm{\xi}_{i}(t-\tau_{ji})-\bm{\xi}_{i})^{2}+\frac{1}{2}(\bm{\xi}_{i}-\bm{\xi}_{i}(t-\tau))^{2}$
$\quad-\frac{1}{2}\bm{\xi}_{i}^{2}(t-\tau)-\frac{1}{2}\bm{\xi}_{i}^{2}\Big{]}$
$\geq\frac{1}{2\tau}\bm{\xi}_{i}^{2}(t-\tau_{ji})-\frac{1}{2\tau}\bm{\xi}_{i}^{2}(t-\tau)-\frac{1}{2\tau}\bm{\xi}_{i}^{2}$
@@
$\dot{\bm{\xi}}_{i}^{2}$ $=[-\kappa\sum_{j=1}^{n}{a_{ij}}{\bm{\xi}}_{i}+{\kappa}\beta\sum_{j=1}^{n}{a_{ij}\bm{\xi}_{j}(t-\tau_{ij})}]^{2}$ (25)
$\leq 2[\sum_{j=1}^{n}{a_{ij}}\kappa{\bm{\xi}}_{i}]^{2}+2\beta^{2}[\sum_{j=1}^{n}a_{ij}\kappa\bm{\xi}_{j}(t-\tau_{ij})]^{2}$
$\leq 2n\sum_{j=1}^{n}{a_{ij}}\kappa^{2}\bm{\xi}_{i}^{2}+2\beta^{2}n\sum_{j=1}^{n}a_{ij}\kappa^{2}\bm{\xi}_{j}^{2}(t-\tau_{ij})$
@@
$\leq 2\gamma_{3}\tau n^{2}\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}\kappa^{2}\bm{\xi}_{i}^{2}+\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ji}\frac{\gamma_{3}}{2\tau}\bm{\xi}_{i}^{2}(t-\tau)$
$\quad+2\gamma_{3}\tau n^{2}\beta^{2}\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}\kappa^{2}\bm{\xi}_{j}^{2}(t-\tau_{ij})$
$\quad-\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ij}\frac{\gamma_{3}}{2\tau}\bm{\xi}_{j}^{2}(t-\tau_{ij})+\sum_{i=1}^{n}\frac{\gamma_{3}}{2\tau}n\bm{\xi}_{i}^{2}$
$\leq\sum_{i=1}^{n}(2\gamma_{3}\tau n^{3}\kappa^{2}+\frac{\gamma_{3}}{2\tau}nI_{6})\bm{\xi}_{i}^{2}+\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ji}\frac{\gamma_{3}}{2\tau}\bm{\xi}_{i}^{2}(t-\tau)$
$\quad+\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}(2\gamma_{3}\tau n^{2}\beta^{2}\kappa^{2}-\frac{\gamma_{3}}{2\tau}I_{6})\bm{\xi}_{j}^{2}(t-\tau_{ij})$
Integrating Eqs.(20)(21)(26) obtains
@@
$\quad-\gamma_{1}\beta\kappa\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ij}(\bm{\xi}_{i}-\frac{1}{2}\bm{\xi}_{j}(t-\tau_{ij}))^{2}$
$\begin{cases}\alpha_{1}=2\gamma_{1}\kappa-2\gamma_{3}\tau n^{3}\kappa^{2}-(\gamma_{2}+\frac{\gamma_{3}}{2\tau}+\gamma_{1}\beta\kappa)nI_{6}\\ \\ \alpha_{2}=(\frac{\gamma_{3}}{2\tau}-\gamma_{1}\beta\kappa)I_{6}-2\gamma_{3}\tau n^{2}\beta^{2}\kappa^{2}\\ \\ \alpha_{3}=(\gamma_{2}-\frac{\gamma_{3}}{2\tau})I_{6}\end{cases}$ (28)
In order to ensure $\dot{V}\leq 0$, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ must be semi-positive definite matrices, which denotes $\alpha_{1},\alpha_{2},\alpha_{3}>0$ since they are diagonal matrices. For a system, the number of robots $n$ and communication delay $\tau$ are determined. The control law is designed by finding the coefficients $k,\beta>0$ such that there exists $\gamma_{1},\gamma_{2},\gamma_{3}>0$ that satisfies $\alpha_{1},\alpha_{2},\alpha_{3}>0$. A simple design criterion is that the coefficients $k,\beta$ are related to the communication delay. When the delay is large, the control coefficients should be smaller. Once the appropriate coefficients are found, it can be concluded that $V\geq 0$ and $\dot{V}\leq 0$. $\dot{V}=0$ if and only if $\bm{\xi}_{i}=0$ for $i=1,\dots,n$. Thus, it can be deduced that the joint error $\bm{\xi}_{i}$ is asymptotically stable according to the Lyapunov theorem. In other words, $\bm{\xi}_{i}\to 0$ as $t\to+\infty$. Then, $\bm{e}_{i}^{\text{d}}-\bm{e}_{i}\to 0$ and $\bm{h}_{i}-\bm{h}_{i}^{\text{ref}}\to 0$. By control law Eq.(9), the interaction wrench between robots and the object can be maintained at the desired target.
@@
Remark 2: The coefficients of Lyapunov function (18) are not specified and the conclusion Eqs.(28) may not seem very intuitive. However, if there is an accurate estimate of the object’s stiffness and then $K^{-1}_{0}K\approx I_{6}$, the coefficients are relatively easy to find in practice because they are all linear inequalities.
III-C Simulations
Figure 2: The interaction wrenches between robots and the object. (a) Without the proposed control law. (b) With the proposed control law.
Figure 3: The communication topology for multiple robots.

### v3

@@ 3.1 Problem Statement and Control Law
Assumption 2: In the control process, the manipulator avoids encountering singular configurations.
@@
Assumption 1 is a very general assumption, which can also be found in [22, 29, 33, 34, 35]. Assumption 2 can be also found in [13, 18, 25]. Assumption 3 is related to the stability proof and will be further discussed in subsequent theoretical proofs.
This paper considers the addition of a compensatory motion to the original motions to achieve the desired interaction wrenches. For the sake of brevity, $\bm{q}(t)$ is abbreviated as $\bm{q}$ in the absence of time delay. For each robot, assuming that $\bm{h}_{i}^{\text{ref}}$ and $\bm{q}_{\text{mm},i}^{\text{ref}}$ have been planned. In planning, the desired interaction wrenches are achieved if $\dot{\bm{q}}_{\text{mm},i}^{\text{ref}}$ is the input. However, due to model deviations and disturbances, this ideal situation is almost impossible. If global positioning and velocity information could be obtained, the control process would be relatively easy since the actual velocity is corrected to match the desired velocity. However, this often requires high-precision and large-scale positioning systems, which are often difficult to implement [28]. Therefore, information from local sensors is considered to design the control laws. Besides, it is also important to consider the communication delay between robots as it is a very common situation in reality. Let $\tau_{ij}(t)$ denote the time delay in the $i$-th robot receiving information from the $j$-th robot. In this paper, the most general delays are considered. The communication delays are non-uniform, asymmetric, time-varying, and unknown. For example, $\tau_{12}(t)\neq\tau_{21}(t)$, $\tau_{12}(t)\neq\tau_{13}(t)$, and thus robots cannot achieve synchronization by querying their own historical information when receiving unknown communication delays.
The actual force exerted on the $i$-th robot during motion by the object is
@@
where $K_{6\times 6}>0$ is a diagonal matrix, representing the stiffness of the object or external environment [8, 36, 29]. Eq.(7) can be rewritten as
@@
where $\bm{e}_{i}^{\text{d}}$ is the desired correction target. Through the force-torque sensor, the direction of correction can be obtained.
In the control of mobile manipulators, redundancy may be introduced. It can be considered when there are additional tasks, such as obstacle avoidance [9, 16, 18], which is not the focus of this paper. Since the range of control correction is generally small and Assumption 2 holds, therefore, only the motions of manipulators are used for correction in the control process [19, 20, 29]. Moreover, the accuracy of the manipulator is higher than the platform, making it worth prioritizing the motion of the manipulator [37]. Considering $\dot{\bm{q}}_{i}=\bm{u}_{i}$ and the control law $\bm{u}_{i}$ is designed as follows
$\bm{u}_{i}=-k$ $\sum_{j=1}^{n}a_{ij}\Big{[}f^{-1}(K^{-1}(\bm{h}_{i}^{\text{ref}}-\bm{h}_{i}))$ (9)
$-\beta f^{-1}(K^{-1}(\bm{h}_{j}^{\text{ref}}(t-\tau_{ij})-\bm{h}_{j}(t-\tau_{ij})))\Big{]}$
where $k$ and $\beta$ are positive coefficients, and $f^{-1}(\cdot)$ is the inverse kinematics of the manipulator. Eq.(9) computes the joint motion of the manipulator to reduce interaction wrenches. The information required for this control law is the local sensor information, the current joint position of the manipulator, and the information from neighboring robots. The control law ensures that the end-effector of the robot moves towards the desired target $\bm{e}_{i}\to\bm{e}_{i}^{\text{d}}$, thereby maintaining the interaction wrench at the expected value $\bm{h}_{i}\to\bm{h}_{i}^{\text{ref}}$. The control framework is shown in Fig. 2, which details how one robot performs feedback control based on the required information. Since each robot individually computes its own motions, without the need for a central controller to compute the motions for all robots, the control law is distributed.
Figure 2: The overall framework. The control process of the $i$-th robot is illustrated.
@@
where $\bm{q}_{i}^{\text{d}}$ is the desired correction joint target. Substituting Eq.(10) into Eq.(9), the control law can be rewritten as
$\bm{u}_{i}=-k\sum_{j=1}^{n}{a_{ij}((\bm{q}_{i}-\bm{q}_{i}^{\text{d}})-\beta(\bm{q}_{j}(t-\tau_{ij})-\bm{q}_{j}^{\text{d}}(t-\tau_{ij})))}$ (11)
Eq.(9) is the control law used in actual calculations, indicating the input and output. The control law is written as Eq.(11) to conveniently illustrate subsequent proofs of stability.
3.2 Stability Analysis
@@
The joint position errors of the $i$-th robot are
@@
$\dot{\bm{\xi}}_{i}=-k\sum_{j=1}^{n}{a_{ij}}{\bm{\xi}}_{i}+k\beta\sum_{j=1}^{n}{a_{ij}\bm{\xi}_{j}(t-\tau_{ij})}$ (17)
For the sake of brevity, the quadratic type $\bm{\xi}_{i}^{\text{T}}\bm{\xi}_{i}$ is abbreviated as $\bm{\xi}_{i}^{2}$. Considering the following Lyapunov function
$V(t)$ $=V_{1}+V_{2}+V_{3}$ (18)
@@
where $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ are the positive coefficients of the Lyapunov function. The Lyapunov function is an equation based on the joint position errors and is positive definite. Taking the time derivative of the Lyapunov function obtains
@@
Substituting Eq.(17) into $\dot{V}_{1}$ obtains
$\dot{V}_{1}\leq-2\gamma_{1}k\sum_{i=1}^{n}\bm{\xi}_{i}^{2}+2\gamma_{1}\beta k\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}\bm{\xi}_{i}^{\text{T}}\bm{\xi}_{j}(t-\tau_{ij})$ (20)
@@
$\dot{V}_{2}\leq\gamma_{2}\sum_{i=1}^{n}n\bm{\xi}_{i}^{2}-\gamma_{2}\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ji}\bm{\xi}_{i}^{2}(t-\tau)$ (21)
@@
$\int_{t-\tau}^{t}{\dot{\bm{\xi}}_{i}^{2}(s)}\text{d}s\geq\frac{1}{2\tau}\bm{\xi}_{i}^{2}(t-\tau_{ji})-\frac{1}{2\tau}\bm{\xi}_{i}^{2}(t-\tau)-\frac{1}{2\tau}\bm{\xi}_{i}^{2}$ (24)
@@
$\dot{\bm{\xi}}_{i}^{2}\leq 2n\sum_{j=1}^{n}{a_{ij}}k^{2}\bm{\xi}_{i}^{2}+2\beta^{2}n\sum_{j=1}^{n}a_{ij}k^{2}\bm{\xi}_{j}^{2}(t-\tau_{ij})$ (25)
@@
$\dot{V}_{3}$ $\leq\sum_{i=1}^{n}\sum_{j=1}^{n}{a_{ij}}(2\gamma_{3}\tau n^{2}\beta^{2}k^{2}-\frac{\gamma_{3}}{2\tau}I_{6})\bm{\xi}_{j}^{2}(t-\tau_{ij})$ (26)
$+\sum_{i=1}^{n}(2\gamma_{3}\tau n^{3}k^{2}+\frac{\gamma_{3}}{2\tau}nI_{6})\bm{\xi}_{i}^{2}$
$+\sum_{i=1}^{n}\sum_{j=1}^{n}a_{ji}\frac{\gamma_{3}}{2\tau}\bm{\xi}_{i}^{2}(t-\tau)$
Integrating Eqs.(20)(21)(26) obtains a negative definite quadratic form
@@
$\begin{cases}\alpha_{1}=2\gamma_{1}k-2\gamma_{3}\tau n^{3}k^{2}-(\gamma_{2}+\frac{\gamma_{3}}{2\tau}+\gamma_{1}\beta k)nI_{6}\\ \\ \alpha_{2}=(\frac{\gamma_{3}}{2\tau}-\gamma_{1}\beta k)I_{6}-2\gamma_{3}\tau n^{2}\beta^{2}k^{2}\\ \\ \alpha_{3}=(\gamma_{2}-\frac{\gamma_{3}}{2\tau})I_{6}\end{cases}$ (28)
In order to ensure $\dot{V}\leq 0$, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ must be semi-positive definite matrices, which denotes $\alpha_{1},\alpha_{2},\alpha_{3}>0$ since they are diagonal matrices. For a system, the number of robots $n$ and communication delay bound $\tau$ are determined. The control law is designed by finding the coefficients $k,\beta>0$ such that there exists $\gamma_{1},\gamma_{2},\gamma_{3}>0$ that satisfies $\alpha_{1},\alpha_{2},\alpha_{3}>0$. A simple design criterion is that the coefficients $k,\beta$ are related to the communication delay bound. When the bound is large, the control coefficients should be smaller. Once the appropriate coefficients are found, it can be concluded that $V\geq 0$ and $\dot{V}\leq 0$. $\dot{V}=0$ if and only if $\bm{\xi}_{i}=0$ for $i=1,\dots,n$.
By using control law Eq.(9), the joint position errors will converge to zero, which in turn drives the end-effector pose errors $\bm{e}_{i}^{\text{d}}-\bm{e}_{i}$ to zero. Furthermore, the interaction wrenches can approach the desired values according to Eq.(10). Therefore, the unnecessary interaction wrenches can be reduced.
@@
3.3 Simulations
Figure 3: The communication topology for multiple robots. (a) Weakly connected graph. (b) Complete graph.